OpenStudy (anonymous):

A large diamond with a mass of 4289.6 grams was recently discovered in a mine. If the density of the diamond is 3.51 grams over centimeters cubed, what is the volume? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

@robtobey @texaschic101

OpenStudy (anonymous):

Density=mass/volume. Volume=mass/density= 4289.6/3.51=1222.1, approximately 1200 to the nearest hundreth.

OpenStudy (anonymous):

Use dimensional analysis.\[\frac{g}{\frac{g}{\text{cm}^3}}=\text{cm}^3 \]

OpenStudy (anonymous):

so wait ?

OpenStudy (anonymous):

Is my answer correct?